Home‎ > ‎Reports‎ > ‎Crew104‎ > ‎

day12


Crew 104 Log Book for May 4, 2011

Photos of the day





Commander's Check-In Report

Jon Clarke Reporting

Crew Physical Status: Good, crew member with a cold is improving.

Time Departed/Returned from EVA:  Three EVAs, 10 am-12, 2 pm-4 pm and 5:30 pm-6:45 pm.

Brief Narrative of Field Mission Results:  all EVAs were to collect final round of clean and sterile samples from previously documented sites.  First was to a point between the Hab and Kissing Camel Ridge, the second was to Giles, and the third to a point just east of the Hab.  Two short EVAs were also done in the morning and afternoon to capture simulated rover images of the two drilling sites.

EVA Data/Interpretations: Pending analysis

Engineering/Hab Maintenance: no outstanding issues

Report Transmission Schedule: commander's and engineering reports later.

Plans for Tomorrow:  Cleaning up, packing, and shipping out equipment to Green River

Inventory: no issues

Support Requested: none

Miscellaneous: none


Commander's Report

Jon Clarke Reporting

May the fourth be with you!

Today is our last field day, and we have been running around like crazy collecting final samples and finishing as many analyses as possible.

The high point for for me today was going to the site of the former town of Giles, and crossing the Fremont River to our sampling site at Steamboat point.  Jhony and I, on Carol's instructions, wanted to collect a clean and sterile sample of the carbonaceous Tununk Shale.  This proved to be 60 meters up a talus slope and at the base of a further 60 m high cliff, the base of which we chipped away nervously. 

We then traversed across the talus slope looking for large carbonate concretions that had been seen previously, to no avail.  We did observe, in the teeth of flying grit, wind ripples and nebkha dunes forming on a 40 degree slope and watched dark slope streaks forming and the sand became unstable and slid down slope, exposing darker underlying talus.  These became lighter in color in a matter of minutes as they were buried by drifting sand.  This may be mimicking the processes that form and obliterate at least some of the Martian slope steaks.


Engineering Report

David Willson Reporting

Generator/Electricity:

Kitty (Used? Yes/No): Used Continuously and not stopped

Honey (Diesel Generator) (Used? Yes/No): Not Used

Wendy (Gas Generator) (Used? Yes/No): Not Used

Battery Charge Level: Not Checked

Notes/Comments (include how many generator runs, approximately when, and any times you needed to turn the generator on/off):

Generator operating very well continuously and not stopped


Fuel Status:

Diesel (Full/0.75/0.5/0.25/Empty): 20% in Tank. The gauge has been reading approx 1/4 - 20% for days.

Propane Tank (Full/0.75/0.5/0.25/Empty): 30%

Gasoline Tank (Full/0.75/0.5/0.25/Empty): N/A

Notes/Comments:

The gauge has been reading approx 1/4 full for days.


Oil:

Kitty Oil Quantity (# of Quarts in storage): Full

Honey Oil Quantity (# of Quarts in storage): Still requires more oil

ATV Oil Quantity (# of Quarts in storage): Not Checked

Car Oil Quantity (# of Quarts in storage): Not Checked


Water Status:

Outside Potable Water Tank Level (inches from bottom): 75 % full

Trailer Potable Water Tank Level (inches from bottom): 12 inches

Hab Potable Water Tank Level (inches from bottom): 16%

Potable Water Meter Reading: 461196

In to GreenHab Meter Reading: N/A

Notes/Comments: Green Hab not operational but water is being recycled. Green Hab tanks are losing water as they are not replenished


GreenHab:

Condition of plants in Tank 1: 5%

Condition of plants in Tank 2: Full

Condition of plants in Tank 3: 15% Full

Greenhab Notes/Comments: Green house not used


Observatory:

Telescope Used? (Yes/No): NO

Observatory Notes/Comments:


Transportation:

Hab Car (NOT RENTAL): Operating

Used (yes/no): NO

Oil Checked (yes/no):

Tire Status:

Odometer at end of day:

Notes/Comments on Hab Car:


ATV 1 (Opportunity): Operating

Used: y/n Yes

Oil Checked: y/n Yes

Fuel Consumed: 20 min

Tire Status: Pumped

Comments: The motor is over-reving & replaced pull cord


ATV 2 (Spirit):

Used: y/n Yes

Oil Checked: y/n Yes

Fuel Consumed: 20 minutes

Tire Status: Pumped

Comments: The auto transmission has trouble changing gear.


ATV 3 (Viking I) (Blue one):

Used: y/n NO

Oil Checked: y/n

Fuel Consumed:

Tire Status:

Comments:


ATV 4 (Viking II):

Used: y/n No

Oil Checked: y/n OK

Fuel Consumed: Full

Tire Status: Pumped

Comments: Not operational. both battery links lost. Motor has trouble starting. Cannot get spark plug out.


Heating and Ventilation:

Thermostat setting upstairs (Fahrenheit): A/C not operating 75 degrees F

Actual temperature upstairs (Fahrenheit): 75 degrees F

Thermostat setting downstairs (Fahrenheit): 73 degrees F

Actual temperature downstairs (Fahrenheit):


Computers/Networking Infrastructure: Operating OK

Notes/comments on internet/computers?: All OK


General Engineering Notes/Comments: